OpenStudy (nuccioreggie):

Simplify twenty-six s plus fourteen t minus nine t minus sixteen s plus nine minus ten s. five t minus ten s plus nine five s t plus nine fourteen t minus nine t five t plus nine

OpenStudy (texaschic101):

26s + 14t - 9t - 16s + 9 - 10s combine like terms.. 26s - 16s - 10s = 10s - 10s = 0 14t - 9t = 5t 9 has no other like term so we are left with : 5t + 9